OpenStudy (anonymous):

What is the equation in standard form which passss through (3, -6) and (-2,-1)

OpenStudy (anonymous):

When two points are given like \((x_1, y_1)\) and \((x_2, y_2)\), then use two point slope form: \[\large \color{blue}{\frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1}}\]

OpenStudy (anonymous):

Which gives 1/-7

OpenStudy (anonymous):

Here : \(x_1 = 3\), \(x_2 = -2\) \(y_1 = -6\). \(y_2 = -1\) Plug in the values and find the equation..

OpenStudy (anonymous):

You cannot substitute for x and y as they are just arbitrary constants.. Your final answer will consist of x and y..

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

Can you solve this: \[\frac{y + 6}{-1 + 6} = \frac{x - 3}{-2-3}\] Cross Multiply and solve..

Directrix (directrix):

@londoneaves22 The slope of the line determined by the points (3, -6) and (-2,-1) is -1. Using the point-slope general form of a line: y - y1 = m(x - x1) where m represents slope and (x1,y1) is a point on the line. I chose (3, -6) which is one of the two points on the line. y - (-6) = (-1)* (x - 3) y + 6 = -1*x + 3 y = -1*x - 3 --> This is the equation of the line in slope-intercept form. What is the equation in standard form? ------------------------ So, you have some work to do to get the equation in standard form. The Standard Form for a linear equation in two variables, x and y, is usually given as Ax + By = C where, if at all possible, A, B, and C are integers, and A is non-negative, and, A, B, and C have no common factors other than 1. y = -1*x - 3 needs to be placed in this form: Ax + By = C.